umankant biswal sir why are you not answering its so urgent I sleep for 2 hours and school timing is 9:00 am - 2:40 pm

Grade:10

1 Answers

Umakant biswal
7 years ago
dear jahanavi 
i am sorry , it takes some time as other student also have questions , , but rest assure i will answer all ques , 
1 thing i would like to say – kindly donot loose sleep at this time , so, at least 6 hours u have to sleep , u are in class 8th now , so, your brain require rest to function more efficiently . 
follow the below time table . 
5;00 – wake up 
5;00-6;00 = meditation , and a bit exercise 
6;05 = 7;00 = daily morning work , 
7;05 = 8:05 = finish the school homework and if no homework then just keep a look what u are gng to study in school . 
8;10 = 8;30 = morning breakfast 
8;40-9;00 – ready for school 
9;00 – 3;05 = school (travelling time included ) 
3;10 – 3;40 = mid day meal ( if u can reach home earlier u can take mid day meal at 2;40 also ) 
3;45 = 4;30 = playing outdoor activity 
4;35 – 6;00 = maths ( 1 hour numerical and half an hour u have to learn the formula ) 
6;05 – 7;05 = science ( alternate in everyday , today , phy , then tomorrow biology then chemi and so on ) 
7;05 = 7;20 = break 
7;30 = 8;30 = social studies . 
8;30 = 8;40 = break 
8;40 = 9;40 = do the past year competetion paper , like nso , imo ,and various ques on ref books , u can prefer HOTS question (high order thinking skill ) 
9;45 = 10;35= language subject ( hindi/sanskrit ) 
10; 40 – 11;05 = night meal 
11;05 = 11;35 = english 
11;35 = go to sleep ,, 
HOPE U CAN FOLLOW IT , IN CASE OF ANY CONCERN LET ME KNOW , 
REGARDS